# SU(3) gauge field of magnons in antiferromagnetic skyrmion crystals

### v2

@@ Abstract
Discussion.—We have shown that the AFM-SkXs give rise to the SU(3) gauge field of magnons, and its commutation relation leads to the uniform flux. The emergence of the uniform flux is particularly important since otherwise the no-go condition precludes a finite thermal Hall conductivity on the triangular lattice [12, 18]. Here, we briefly demonstrate this fact by considering the triangular-lattice ferromagnet with the DM interaction shown in Fig. 3(a). We assume that the direction of the spin in the classical ground state aligns with that of the DM vector $\bm{D}_{i,j}$. The DM interaction leads to the complex hopping of magnons as $J\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}+\bm{D}_{i,j}\cdot(\hat{\bm{S}}_{i}\times\hat{\bm{S}}_{j})\simeq\sqrt{J^{2}+D^{2}}S(\mathrm{e}^{i\phi_{\mathrm{D}}/3}\hat{b}_{i}^{\dagger}\hat{b}_{j}+\mathrm{H.c.})$, where $D=|\bm{D}_{i,j}|$, $\phi_{\mathrm{D}}/3=-\arctan(D/J)$, and $\hat{b}_{i}$ ($\hat{b}_{i}^{\dagger}$) is the magnon annihilation (creation) operator at site $i$. The DM-induced complex hopping results in the staggered flux pattern shown in Fig. 3(b), which is preserved under a $\pi$-rotation along any straight line in the triangular lattice. This rotation, however, converts the sign of the thermal Hall conductivity, indicating that the DM-induced flux does not contribute to the magnon thermal Hall effect. The FM-SkXs and AFM-SkXs, in contrast, generate the uniform flux patterns due to the finite skyrmion density, thereby circumventing the no-go condition and contributing to the magnon thermal Hall effect.
Figure 3: Staggered flux in the triangular-lattice ferromagnet with the DM interaction. (a) Schematic illustration of the triangular-lattice ferromagnet with the DM interaction. Gray arrows indicate the order of spins in $\bm{D}_{i,j}\cdot(\hat{\bm{S}}_{i}\times\hat{\bm{S}}_{j})$. The magnons acquire a phase $\phi_{\mathrm{D}}/3=-\arctan(D/J)$ when they hop between two sites. (b) Staggered flux pattern generated by the DM interaction. The system has the $\pi$-rotation symmetry along any straight line in the triangular lattice, which prohibits the finite thermal Hall conductivity.
@@
Conclusions and outlook.—We have shown that low-energy excitations in the AFM-SkXs on the triangular lattice are effectively governed by the magnons with the SU(3) gauge field. We have also demonstrated that the finite commutation relation of the SU(3) gauge field breaks the effective time-reversal symmetry and underpins a measurable magnon thermal Hall effect. There are several promising directions for future work. For example, it would be interesting to explore fictitious gauge fields in different classes of skyrmions [48], including two-sublattice AFM-SkXs [49, 50, 51]. Our framework is readily applicable to these systems. Elucidating how an underlying gauge structure qualitatively changes a thermal Hall conductivity would also offer an interesting research direction. Moreover, we could verify the effective field-theoretical description by comparing magnon bands and physical quantities with those obtained by the linear spin-wave theory.

### v3

@@ Abstract
Thermal Hall conductivity.—To demonstrate that the finite field strength from the AFM-SkX spin texture is responsible for the magnon thermal Hall effect, we calculate the thermal Hall conductivity by the linear spin-wave theory. We first determine the spin configurations in the classical ground state by the mean-field approximation. Applying the Holstein-Primakoff transformation, the spin model (1) is reduced to the quadratic magnon Hamiltonian $\hat{\mathcal{H}}_{\mathrm{mag}}=(1/2)\sum_{\bm{k}}\hat{\Phi}_{\bm{k}}^{\dagger}H_{\mathrm{BdG}}(\bm{k})\hat{\Phi}_{\bm{k}}$ up to the constant term, where $\hat{\Phi}_{\bm{k}}=(\hat{b}_{\bm{k},1},\cdots,\hat{b}_{\bm{k},N_{\mathrm{m}}},\hat{b}_{-\bm{k},1}^{\dagger},\cdots,\hat{b}_{-\bm{k},N_{\mathrm{m}}}^{\dagger})^{T}$, $\hat{b}_{\bm{k},s}$ ($\hat{b}_{\bm{k},s}$) is the magnon annihilation (creation) operator with wave vector $\bm{k}$ and magnetic sublattice index $s$, $N_{\mathrm{m}}$ is the number of sites in a magnetic unit cell, and $H_{\mathrm{BdG}}(\bm{k})$ is the $2N_{\mathrm{m}}\times 2N_{\mathrm{m}}$ Hermitian matrix containing systems’ details. The Bogoliubov transformation, which preserves the bosonic commutation relations, introduces new bosonic operators $\hat{\Phi}_{\bm{k}}=P(\bm{k})\hat{\Psi}_{\bm{k}}$, where $\hat{\Psi}_{\bm{k}}=(\hat{\beta}_{\bm{k},1},\cdots,\hat{\beta}_{\bm{k},N_{\mathrm{m}}},\hat{\beta}_{-\bm{k},1}^{\dagger},\cdots,\hat{\beta}_{-\bm{k},N_{\mathrm{m}}}^{\dagger})^{T}$ and $P(\bm{k})$ is the paraunitary matrix that diagonalizes $H_{\mathrm{BdG}}(\bm{k})$ [53]. The magnon Hamiltonian then takes the diagonal form $\hat{\mathcal{H}}_{\mathrm{mag.}}=\sum_{\bm{k}}\sum_{n=1}^{N_{\mathrm{m}}}\varepsilon_{n}(\bm{k})(\hat{\beta}_{\bm{k},n}^{\dagger}\hat{\beta}_{\bm{k},n}+1/2)$ with the $n$th magnon band $\varepsilon_{n}(\bm{k})$. See Supplemental Material for more details [51].
The thermal Hall conductivity in the linear response regime is given by [17, 16, 54]
$\kappa_{xy}=-\frac{k_{\mathrm{B}}^{2}T}{\hbar}\int_{\mathrm{BZ}}\frac{\mathrm{d}^{2}\bm{k}}{(2\pi)^{2}}\sum_{n=1}^{N_{\mathrm{m}}}c_{2}[f(\varepsilon_{n}(\bm{k}))]\Omega_{xy}^{(n)}(\bm{k}),$ (23)
where $c_{2}[x]=\int_{0}^{x}dt[\ln(1+t^{-1})]^{2}$, $f(\varepsilon)=1/[\exp(\varepsilon/k_{\mathrm{B}}T)-1]$ is the Bose distribution function, and $\Omega_{xy}^{(n)}(\bm{k})=-2\mathrm{Im}[\{\partial_{k_{x}}T^{\dagger}(\bm{k})\}\Sigma^{z}\{\partial_{k_{y}}T(\bm{k})\}]_{n,n}$ is the Berry curvature of $n$th magnon bands, and $\Sigma^{z}=\tau^{z}\otimes I_{N_{\mathrm{m}}\times N_{\mathrm{m}}}$ with $N_{\mathrm{m}}\times N_{\mathrm{m}}$ unit matrix $I_{N_{\mathrm{m}}\times N_{\mathrm{m}}}$. Figures 3(a) and 3(b) show the magnetic-field dependence of the thermal Hall conductivity with (a) JS=1.0JS=-1.0 and $D/|J|=1.0$ and (b) $JS=1.0$ and $D/|J|=0.5$  [^1] We choose different values of $D/|J|$ for the FM-SkX and AFM-SkX so that the spin configurations fit within a numerically tractable magnetic unit cell. See Supplemental Material for details [51].. In both the FM and AFM cases, $\kappa_{xy}$ remains very small in the helical phases and becomes very large upon entering the skyrmion phases. This indicates that FM-SkX (AFM-SkX) spin texture is the primary origin of the magnon thermal Hall effect, supporting our field-theoretical results.
Finally, we discuss how the distinct gauge structures of the FM-SkXs and AFM-SkXs possibly influence the thermal Hall conductivity. While field-theoretical approaches often lack quantitative precision, they may still capture the overall tendency in the thermal Hall conductivity. In the FM-SkXs, the finite skyrmion density produces the uniform component of the effective magnetic field and leads to magnon Landau Levels, whose spacing is proportional to the field [20, 28]. A large skyrmion density then pushes most Landau levels out of the low-energy region. Since magnons obey Bose statistics, the thermal Hall conductivity decreases with increasing the skyrmion density, in contrast to electronic systems [26]. In the AFM-SkXs, the amplitude of the field strength from the spin texture is also proportional to the skyrmion density of the vector field $\bm{n}(\bm{r})$, whose amplitude is identical to the skyrmion density of the FM-SkX on each sublattice. However, the prefactor $1/8$ and additional terms from the DM interaction in Eq. (22) imply that the effective magnetic field in the AFM-SkXs tends to be smaller than that in the FM-SkXs. Therefore, we roughly expect that the thermal Hall conductivity of the AFM-SkXs will be larger than that of the FM-SkXs.
Figure 3: Magnon bands and thermal Hall conductivity. (a), (b) Magnetic-field dependence of the thermal Hall conductivity at $k_{\mathrm{B}}T=0.1,0.2,0.3$ for (a) $J<0$ and (b) $J>0$. (c), (d) Magnon bands along the symmetric line in the Brillouin zone for the (c) FM-SkX phase with $g\mu_{\mathrm{B}}B=0.3$ and (d) AFM-SkX phase with $g\mu_{\mathrm{B}}B=3.0$. The color density represents the Berry curvature. We set $|J|S=1.0$ and $D/|J|=1.0$ ($0.5$) for $J<0$ ($J>0$).
Figures 3(c) and 3(d) show the magnon bands and Berry curvatures in the FM-SkX phase with JS=1.0JS=-1.0, $D/|J|=1.0$, and $g\mu_{\mathrm{B}}B=0.3$ and AFM-SkX phase with $JS=1.0$, $D/|J|=0.5$, and $g\mu_{\mathrm{B}}B=3.0$ along the symmetric line in the Brillouin zone. Here, $\Gamma$, $K$, and $M$ points in Figs. 3(c) and 3(d) correspond to $\bm{k}=(0.0)$, $(4\pi/3L_{\mathrm{m}},0)$, and $(2\pi/3L_{\mathrm{m}},\pi/\sqrt{3}L_{\mathrm{m}})$, where $L_{\mathrm{m}}=9$ (77) for the FM-SkX (AFM-SkX) and $N_{\mathrm{m}}=L_{\mathrm{m}}^{2}$. The amplitude of the skyrmion density is numerically estimated as $|\rho_{\mathrm{FM-SkX}}|\simeq 8.9\times 10^{-3}$ for the FM-SkX phase with $g\mu_{\mathrm{B}}B=0.3$ and $\rho_{\mathrm{AFM-SkX}}\simeq 1.2\times 10^{-2}$ for the AFM-SkX phase with $g\mu_{\mathrm{B}}=3.0$ [51]. Despite $|\rho_{\mathrm{AFM-SkX}}|>|\rho_{\mathrm{FM-SkX}}|$, the magnon bands in the AFM-SkX are narrower and more densely spaced than those in the FM-SkX, suggesting a larger thermal Hall conductivity in the AFM-SkX. The thermal Hall conductivity in the AFM-SkX phase (Fig. (3(b))) is indeed larger than that in the FM-SkX (Fig. 3(a)), supporting our conjecture. A more quantitative understanding of the relationship between the field strength and thermal Hall conductivity in AFM-SkXs phase remains for future work.
Conclusions and outlook.—We have shown that low-energy excitations in the AFM-SkXs on the triangular lattice are effectively governed by the magnons with the SU(3) gauge field. We have also demonstrated that the finite commutation relation of the SU(3) gauge field breaks the effective time-reversal symmetry and underpins a measurable magnon thermal Hall effect. The emergence of the thermal Hall conductivity in the AFM-SkX phase has also been demonstrated by the linear spin-wave theory. The relation between the gauge structures and magnitude of the thermal Hall conductivity has also been discussed. In contrast to earlier field-theoretical approaches that rely on the Lagrangian formalism and involve integrating out uniform magnetization components [56, 57], our formulation is simpler and proceeds in a straightforward manner, analogous to the standard Holstein-Primakoff transformation. We have explicitly derived the connection between the AFM-SkX spin texture and the SU(3) gauge field, and we have further clarified how its non-commutative nature leads to the effective time-reversal symmetry breaking.
@@
There are several promising directions for future work. For example, it would be interesting to explore fictitious gauge fields in different classes of skyrmions [58], including two-sublattice AFM-SkXs [59, 60, 61]. Our framework is readily applicable to these systems. The detailed analysis of the relation between gauge structures and thermal Hall conductivity is left for future research. Elucidating how an underlying gauge structure qualitatively changes other physical quantities would also offer an interesting research direction.